For the year ended December 31, 2018, Norstar Industries reported net income of $945,000. At January 1, 2018, the company had 880,000 common shares outstanding. The following changes in the number of shares occurred during 2018:

Apr. 30 Sold 65,000 shares in a public offering.
May 24 Declared and distributed a 5% stock dividend.
June 1 Issued 66,000 shares as part of the consideration for the purchase of assets from a subsidiary.


Required:

Compute Norstar's earnings per share for the year ended December 31, 2018.
Answer #1

Computation of earnings per share is:

Earnings per share = Net income / Weighted outstanding shares

= $945,000 / 1,008,000

= $0.9375

Hence, the earnings per share is $0.9375

Working Note:

1.

Computation of shares from January 1 to December 31 is:

Number of outstanding shares = Number of shares outstanding at January 1 * (1 + Percentage of stock dividend)

= 880,000 * (1 + 0.05)

= 880,000 * 1.05

= 924,000

Hence, the number of shares outstanding is 924,000.

2.

Computation of shares from April 30 to December 31 is:

Number of outstanding shares = Number of shares sold in public offering * (1 + Percentage of stock dividend) * (Number of months from April 30 to December 31 / Number of months annually)

= 65,000 * (1 + 0.05) * (8/12)

= 65,000 * 1.05 * (8/12)

= 45,500

Hence, the number of outstanding shares is 45,500.

3.

Computation of shares from June 1 to December 31 is:

Number of outstanding shares = Number of shares issued * (Number of months from June 1 to December 31 / Number of months annually)

= 66,000 * (7/12)

= 38,500

Hence, the number of outstanding shares is 38,500.

4.

Computation of weighted outstanding shares is:

Weighted outstanding shares = Shares from January to December + Shares from April 30 to December 31 + Shares from June 1 to December 31

= 924,000 + 45,500 + 38,500

= 1,008,000

Hence, the weighted outstanding shares is 1,008,000.